Compound ‘A’ with the molecular formula C5H8 reacts with hydrogen in the presence of Lindlar’s catalyst to form a compound ‘B’ with the molecular formula C5H10. ‘A’ on reacting with sodium in liquid ammonia forms a compound ‘C’ with same molecular formula as that of ‘B’. Identify A, B and C. Also write the chemical reactions involved.

Grade:12

1 Answers

Sunil Kumar FP
9 years ago
Compound ‘A’ with the molecular formula C5H8 reacts with hydrogen in the presence of Lindlar’s catalyst to form a compound ‘B’ with the molecular formula C5H10. ‘A’ on reacting with sodium in liquid ammonia forms a compound ‘C’ with same molecular formula as that of ‘B’
R1-C≡C-R2 + H2 & Lindlar catalyst ——> cis R1-CH=CH-R2
R1-C≡C-R2 + 2 Na in NH3 (liq) ——> trans R1-CH=CH-R2 + 2 NaNH2
here R1=CH3CH2 and R2=CH3
A=CH3CH2CCCH3 and B=CH3CH2CHCHCH3 and C=>CH3CH2CHCHCH3